# Statistical Inference in Reinforcement Learning: A Selective Survey

### v1

@@ 3.2 Simulations: Tiger problem
Figure 3: Left: The tiger problem. Middle: Percentage of rejections of the null hypothesis that the data from the tiger problem satisfies a $k$-th order Markov assumption using Shi et al. (2020)’s test, for $k=1,2,\cdots,10$. $n$ denotes the number of trajectories. Right: Same as Middle, but with the true location of the tiger known and included in the observations.
Shi et al. (2020) applied their test to this environment to test whether the generated data satisfies a $k$-th order MDP, where $k=1,2,\cdots,10$. The significance level was set to 0.05, and we plot the proportion of rejections of the null hypothesis in 500 random simulations in the middle panel of Figure 3. Since the true model follows a POMDP, the null hypothesis should theoretically be rejected regardless of the value of $k$. Thus, the proportions in the middle panel reflect the power of their test. It can be seen that the power increases as the number of trajectories $n$ grows and is very close to $1$ when $n\geq 100$ and $k\leq 7$. This indicates that the test can effectively distinguish POMDPs from (higher-order) MDPs.
Additionally, Shi et al. (2020) conducted another simulation by including the true location of the tiger in the observation. In this case, the environment satisfies the MDP assumption due to the full observability of the state, and the null hypothesis holds regardless of the value of $k$. We also plot the proportion of rejections of the null hypothesis – which corresponds to the type-I error – in the right panel of Figure 3, with upper and lower horizontal lines depicting the Monte Carlo error bounds. The results show that most proportions lie within the error bounds, demonstrating that the test can effectively identify (high-order) MDPs.
@@ 4 Methodology: Forward-backward (Generative) Learning
Challenge. In this section, we focus on reviewing the hypothesis tests proposed by Shi et al. (2020) and Zhou et al. (2023), as both papers studied the RL setting [^5] Zhou et al. (2023) developed the test in the classic time series setting and later extended it to MDPs in Section 6.. Later, we will discuss other relevant tests designed for conditional independence or the Markov assumption in time series analysis. Before delving into the details, we first outline the main challenges in testing the Markov property in RL and how these methods effectively address them. The primary challenge arises from the presence of high-dimensional continuous conditional variables. As we will discuss later, the Markov assumption is equivalent to a set of conditional independence conditions. Testing conditional independence is considerably more difficult than testing marginal independence when the conditional variables are continuous (Bergsma, 2004). This challenge is further compounded when these conditional variables are high-dimensional. However, high-dimensional continuous variables are common in RL applications. For example, testing a high-order MDP requires to merge the current observation with past lagged variables to define the state, which naturally introduces high-dimensional conditional variables. In the diabetes study, variables such as blood glucose levels and calorie intake are continuous.
@@
Shi et al. (2020) and Zhou et al. (2023) leveraged modern supervised learning and deep generative learning algorithms to handle high-dimensional continuous conditional variables, respectively. Take Zhou et al. (2023)’s proposal as an example. The idea is to learn two generators: (i) a forward generator, which predicts the next observation based on the current observation-action pair, and (ii) a backward generator, which reconstructs the current observation and action based on the next observation-action pair. By combining these two generators, Zhou et al. (2023) constructed a doubly robust test statistic, effectively eliminating the bias from a single generator and ensuring the test statistic has a desirable limiting distribution. Shi et al. (2020) constructed a similar doubly robust test statistic with a forward learner and a backward learner. We elaborate on these forward-backward-learning-based methods below.
@@
Null and alternative hypotheses. To simplify the discussion, we assume that the reward ${\color[rgb]{1,0,0}\bm{R}_{t}}$ is a deterministic function of the observation ${\color[rgb]{1,.5,0}\bm{O}_{t}}$, action ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$, and the next observation ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$. This assumption holds automatically in the diabetes case study, where ${\color[rgb]{1,0,0}\bm{R}_{t}}$ depends only on the blood glucose level at time $t+1$, which is included in ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$. To ensure this assumption is always satisfied, we can redefine the observation by incorporating ${\color[rgb]{1,0,0}\bm{R}_{t}}$ into ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$. Based on this assumption, we consider the following null and alternative hypotheses:
@@
  • $\mathcal{H}_{0}$: For any time $t$, given ${\color[rgb]{1,.5,0}\bm{O}_{t}}$ and ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$, ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$ is conditionally independent of past observation-action pairs $({\color[rgb]{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{i}})_{i<t}$.
@@
  • $\mathcal{H}_{a}$: There exists at least some $t$ where ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$ is not conditionally independent of past observation-action pairs given ${\color[rgb]{1,.5,0}\bm{O}_{t}}$ and ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$.
@@
According to Theorem 1 in Shi et al. (2020) and Zhou et al. (2023), under the stationarity assumption, the null hypothesis is equivalent to the intersection of a series of conditional independence hypotheses $\cap_{q>1}\mathcal{H}_{0}(q)$ where
@@
  • $\mathcal{H}_{0}(q)$: For any time $t$ and integer $q>1$, given $({\color[rgb]{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{i}})_{t<i<t+q}$, ${\color[rgb]{1,.5,0}\bm{O}_{t+q}}$ is conditionally independent of $({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})$.
@@
$\mathbb{E}\{g({\color[rgb]{1,.5,0}\bm{O}_{t+q}})-\mathbb{E}[g({\color[rgb]{1,.5,0}\bm{O}_{t+q}})|({\color[rgb]{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{i}})_{t<i<t+q}]\}\{h({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})-\mathbb{E}[h({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})|({\color[rgb]{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{i}})_{t<i<t+q}]\}=0.$
@@
Using the Markov assumption, it is immediate to see that the above two conditional expectations depend only on $({\color[rgb]{1,.5,0}\bm{O}_{t+q-1}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t+q-1}})$ and $({\color[rgb]{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t+1}})$, respectively. This leads to
$\mathbb{E}\{g({\color[rgb]{1,.5,0}\bm{O}_{t+q}})-\mathbb{E}[g({\color[rgb]{1,.5,0}\bm{O}_{t+q}})|({\color[rgb]{1,.5,0}\bm{O}_{t+q-1}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t+q-1}})]\}\{h({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})-\mathbb{E}[h({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})|{\color[rgb]{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t+1}}]\}=0.$ (1)
@@
Equation (1) is the key to the tests by Shi et al. (2020) and Zhou et al. (2023). Specifically, both estimate the two conditional expectations $\mathbb{E}[g({\color[rgb]{1,.5,0}\bm{O}_{t+1}})|{\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}]$ and $\mathbb{E}[h({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})|{\color[rgb]{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t+1}}]$ for a set of functions $g$ and $h$, plug-in these estimators into Equation (1), and approximate the leftmost expectation using the sample average to construct the test statistic $S(q,g,h)$. Under the null hypothesis, $S(q,g,h)$ should be close to 0 for all $q$, $g$, and $h$. If there exist combinations of $q$, $g$, and $h$ for which the statistic $S(q,g,h)$ significantly deviates from 0, this provides evidence to reject the null hypothesis, indicating the violation of the Markov property. Based on this, Shi et al. (2020) and Zhou et al. (2023) considered a large number of combinations of $q$, $g$, and $h$, and construct a maximum-type test statistic $\max_{q,g,h}|S(q,g,h)|$. If this test statistic exceeds a certain threshold (discussed below), the null is rejected.
@@
  • According to Theorem 2 in Shi et al. (2020) and Zhou et al. (2023), under the null hypothesis, Equation (1) holds as long as either one of the two conditional expectations, $\mathbb{E}[g({\color[rgb]{1,.5,0}\bm{O}_{t+1}})|{\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}]$ or $\mathbb{E}[h({\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}})|{\color[rgb]{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t+1}}]$ is correctly specified. Such a double robustness property greatly facilitates hypothesis testing. Specifically, when the convergence rates of these conditional expectations are $o_{p}(N^{-1/4})$ where $N$ denotes the total sample size (i.e., the total number of observation-action-next-observation triplets), the resulting test becomes consistent. We do not require them to achieve the standard parametric convergence rate; see Condition (C4) in Shi et al. (2020) and Condition 4 in Zhou et al. (2023).
@@
  • Given the need to consider a large number of $g$ and $h$ combinations, it is essential to develop an efficient algorithm for computing the conditional expectations. Shi et al. (2020) proposed to employ random forests to model these conditional expectations and adapted the quantile random forest algorithm (Meinshausen and Ridgeway, 2006) to accelerate computation. Specifically, in their proposal, the structure of the forest (e.g., the tree depth and the split points) is independent of $g$ (or $h$), meaning that the same forest is used for any $g$ (or $h$). The final conditional expectation is computed based on the specific form of $g$ (or $h$), which substantially simplifies the computation since the forest does not need to be retrained for different $g$ (or $h$). For details, see Section 5.1 of Shi et al. (2020).
@@
  • Similarly, Zhou et al. (2023) employ generative learning algorithms to improve the computational efficiency. Take the computation of $\mathbb{E}[g({\color[rgb]{1,.5,0}\bm{O}_{t+1}})|{\color[rgb]{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}]$ as an example. The key idea is to learn the conditional distribution of ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$ given ${\color[rgb]{1,.5,0}\bm{O}_{t}}$ and ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$. This requires to learn a generator that takes ${\color[rgb]{1,.5,0}\bm{O}_{t}}$, ${\color[rgb]{0.0,0.5,0.0}\bm{A}_{t}}$, and random noise as inputs and outputs ${\color[rgb]{1,.5,0}\bm{O}^{*}}$, such that the distribution of ${\color[rgb]{1,.5,0}\bm{O}^{*}}$ closely approximates the conditional distribution of ${\color[rgb]{1,.5,0}\bm{O}_{t+1}}$. By generating $M$ independent samples $\{{\color[rgb]{1,.5,0}\bm{O}^{(m)}}\}_{m=1}^{M}$ from this generator, the conditional expectation can be approximated as $M^{-1}\sum_{m=1}^{M}g({\color[rgb]{1,.5,0}\bm{O}^{(m)}})$. In the experiments, Zhou et al. (2023) used the mixture density network (Bishop, 1994) to model the conditional distribution. Meanwhile, other generative neural network models, such as generative adversarial networks (GANs, Goodfellow et al., 2014) and diffusion models (Ho et al., 2020) are equally applicable; see e.g., Bellot and van der Schaar (2019), Shi et al. (2021, 2023) and Zhang et al. (2024, 2025) for their use in conditional (mean) independence testing.
@@
  • Both Shi et al. (2020) and Zhou et al. (2023) set $g$ and $h$ to characteristic functions because they uniquely determine the distribution function. Alternatively, neural networks can also be used for $g$ and $h$, as suggested by Shi et al. (2021).
@@
Specifically, under the null hypothesis, all $S(q,g,h)$ asymptotically follow a multivariate Gaussian distribution. Thus, $\max_{q,g,h}|S(q,g,h)|$ converges in distribution to the maximum of several multivariate Gaussian variables. We first estimate the covariance matrix of these statistics using their sample covariance. Then, we simulate random vectors from a Gaussian distribution with the same covariance matrix, take the absolute value of each component, and compute the maximum. By repeating this process, we obtain a series of bootstrap statistics. Finally, the threshold is set to the upper $\alpha$-quantile of these bootstrap statistics. If the test statistic exceeds this threshold, the null hypothesis is rejected. Theoretical properties of this test are detailed in Theorem 1 of Shi et al. (2020) and Theorems 5 and 6 of Zhou et al. (2023).

### v2

@@ 3.2 Simulations: Tiger problem
Figure 3: Left: The tiger problem. Middle: Percentage of rejections of the null hypothesis that the data from the tiger problem satisfies a $k$-th order Markov assumption using Shi et al. (2020b)’s test, for $k=1,2,\cdots,10$. $n$ denotes the number of trajectories. Right: Same as Middle, but with the true location of the tiger known and included in the observations.
Shi et al. (2020b) applied their test to this environment to test whether the generated data satisfies a $k$-th order MDP, where $k=1,2,\cdots,10$. The significance level was set to 0.05, and we plot the proportion of rejections of the null hypothesis in 500 random simulations in the middle panel of Figure 3. Since the true model follows a POMDP, the null hypothesis should theoretically be rejected regardless of the value of $k$. Thus, the proportions in the middle panel reflect the power of their test. It can be seen that the power increases as the number of trajectories $n$ grows and is very close to $1$ when $n\geq 100$ and $k\leq 7$. This indicates that the test can effectively distinguish POMDPs from (higher-order) MDPs.
Additionally, Shi et al. (2020b) conducted another simulation by including the true location of the tiger in the observation. In this case, the environment satisfies the MDP assumption due to the full observability of the state, and the null hypothesis holds regardless of the value of $k$. We also plot the proportion of rejections of the null hypothesis – which corresponds to the type-I error – in the right panel of Figure 3, with upper and lower horizontal lines depicting the Monte Carlo error bounds. The results show that most proportions lie within the error bounds, demonstrating that the test can effectively identify (high-order) MDPs.
@@ 4 Methodology: Forward-backward (Generative) Learning
Challenge. In this section, we focus on reviewing the hypothesis tests proposed by Shi et al. (2020b) and Zhou et al. (2023b), as both papers studied the RL setting [^5] Zhou et al. (2023b) developed the test in the classic time series setting and later extended it to MDPs in Section 6.. Later, we will discuss other relevant tests designed for conditional independence or the Markov assumption in time series analysis. Before delving into the details, we first outline the main challenges in testing the Markov property in RL and how these methods effectively address them. The primary challenge arises from the presence of high-dimensional continuous conditional variables. As we will discuss later, the Markov assumption is equivalent to a set of conditional independence conditions. Testing conditional independence is considerably more difficult than testing marginal independence when the conditional variables are continuous (Bergsma, 2004). This challenge is further compounded when these conditional variables are high-dimensional. However, high-dimensional continuous variables are common in RL applications. For example, testing a high-order MDP requires to merge the current observation with past lagged variables to define the state, which naturally introduces high-dimensional conditional variables. In the diabetes study, variables such as blood glucose levels and calorie intake are continuous.
@@
Shi et al. (2020b) and Zhou et al. (2023b) leveraged modern supervised learning and deep generative learning algorithms to handle high-dimensional continuous conditional variables, respectively. Take Zhou et al. (2023b)’s proposal as an example. The idea is to learn two generators: (i) a forward generator, which predicts the next observation based on the current observation-action pair, and (ii) a backward generator, which reconstructs the current observation and action based on the next observation-action pair. By combining these two generators, Zhou et al. (2023b) constructed a doubly robust test statistic, effectively eliminating the bias from a single generator and ensuring the test statistic has a desirable limiting distribution. Shi et al. (2020b) constructed a similar doubly robust test statistic with a forward learner and a backward learner. We elaborate on these forward-backward-learning-based methods below.
@@
Null and alternative hypotheses. To simplify the discussion, we assume that the reward ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}}$ is a deterministic function of the observation ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$, action ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$, and the next observation ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$. This assumption holds automatically in the diabetes case study, where ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}}$ depends only on the blood glucose level at time $t+1$, which is included in ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$. To ensure this assumption is always satisfied, we can redefine the observation by incorporating ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{R}_{t}}$ into ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$. Based on this assumption, we consider the following null and alternative hypotheses:
@@
  • $\mathcal{H}_{0}$: For any time $t$, given ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$ and ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$, ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$ is conditionally independent of past observation-action pairs $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{i}})_{i<t}$.
@@
  • $\mathcal{H}_{a}$: There exists at least some $t$ where ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$ is not conditionally independent of past observation-action pairs given ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$ and ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$.
@@
According to Theorem 1 in Shi et al. (2020b) and Zhou et al. (2023b), under the stationarity assumption, the null hypothesis is equivalent to the intersection of a series of conditional independence hypotheses $\cap_{q>1}\mathcal{H}_{0}(q)$ where
@@
  • $\mathcal{H}_{0}(q)$: For any time $t$ and integer $q>1$, given $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{i}})_{t<i<t+q}$, ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+q}}$ is conditionally independent of $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})$.
@@
$\mathbb{E}\{g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+q}})-\mathbb{E}[g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+q}})|({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{i}})_{t<i<t+q}]\}\{h({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})-\mathbb{E}[h({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})|({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{i}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{i}})_{t<i<t+q}]\}=0.$
@@
Using the Markov assumption, it is immediate to see that the above two conditional expectations depend only on $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+q-1}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t+q-1}})$ and $({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t+1}})$, respectively. This leads to
$\mathbb{E}\{g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+q}})-\mathbb{E}[g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+q}})|({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+q-1}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t+q-1}})]\}\{h({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})-\mathbb{E}[h({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t+1}}]\}=0.$ (1)
@@
Equation (1) is the key to the tests by Shi et al. (2020b) and Zhou et al. (2023b). Specifically, both estimate the two conditional expectations $\mathbb{E}[g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}})|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}]$ and $\mathbb{E}[h({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t+1}}]$ for a set of functions $g$ and $h$, plug-in these estimators into Equation (1), and approximate the leftmost expectation using the sample average to construct the test statistic $S(q,g,h)$. Under the null hypothesis, $S(q,g,h)$ should be close to 0 for all $q$, $g$, and $h$. If there exist combinations of $q$, $g$, and $h$ for which the statistic $S(q,g,h)$ significantly deviates from 0, this provides evidence to reject the null hypothesis, indicating the violation of the Markov property. Based on this, Shi et al. (2020b) and Zhou et al. (2023b) considered a large number of combinations of $q$, $g$, and $h$, and construct a maximum-type test statistic $\max_{q,g,h}|S(q,g,h)|$. If this test statistic exceeds a certain threshold (discussed below), the null is rejected.
@@
  • According to Theorem 2 in Shi et al. (2020b) and Zhou et al. (2023b), under the null hypothesis, Equation (1) holds as long as either one of the two conditional expectations, $\mathbb{E}[g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}})|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}]$ or $\mathbb{E}[h({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}})|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t+1}}]$ is correctly specified. Such a double robustness property greatly facilitates hypothesis testing. Specifically, when the convergence rates of these conditional expectations are $o_{p}(N^{-1/4})$ where $N$ denotes the total sample size (i.e., the total number of observation-action-next-observation triplets), the resulting test becomes consistent. We do not require them to achieve the standard parametric convergence rate; see Condition (C4) in Shi et al. (2020b) and Condition 4 in Zhou et al. (2023b).
@@
  • Given the need to consider a large number of $g$ and $h$ combinations, it is essential to develop an efficient algorithm for computing the conditional expectations. Shi et al. (2020b) proposed to employ random forests to model these conditional expectations and adapted the quantile random forest algorithm (Meinshausen and Ridgeway, 2006) to accelerate computation. Specifically, in their proposal, the structure of the forest (e.g., the tree depth and the split points) is independent of $g$ (or $h$), meaning that the same forest is used for any $g$ (or $h$). The final conditional expectation is computed based on the specific form of $g$ (or $h$), which substantially simplifies the computation since the forest does not need to be retrained for different $g$ (or $h$). For details, see Section 5.1 of Shi et al. (2020b).
@@
  • Similarly, Zhou et al. (2023b) employ generative learning algorithms to improve the computational efficiency. Take the computation of $\mathbb{E}[g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}})|{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}},{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}]$ as an example. The key idea is to learn the conditional distribution of ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$ given ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$ and ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$. This requires to learn a generator that takes ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t}}$, ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{A}_{t}}$, and random noise as inputs and outputs ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}^{*}}$, such that the distribution of ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}^{*}}$ closely approximates the conditional distribution of ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}_{t+1}}$. By generating $M$ independent samples $\{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}^{(m)}}\}_{m=1}^{M}$ from this generator, the conditional expectation can be approximated as $M^{-1}\sum_{m=1}^{M}g({\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\bm{O}^{(m)}})$. In the experiments, Zhou et al. (2023b) used the mixture density network (Bishop, 1994) to model the conditional distribution. Meanwhile, other generative neural network models, such as generative adversarial networks (GANs, Goodfellow et al., 2014) and diffusion models (Ho et al., 2020) are equally applicable; see e.g., Bellot and van der Schaar (2019), Shi et al. (2021b, 2023b) and Zhang et al. (2024, 2025) for their use in conditional (mean) independence testing.
@@
  • Both Shi et al. (2020b) and Zhou et al. (2023b) set $g$ and $h$ to characteristic functions because they uniquely determine the distribution function. Alternatively, neural networks can also be used for $g$ and $h$, as suggested by Shi et al. (2021b).
@@
Specifically, under the null hypothesis, all $S(q,g,h)$ asymptotically follow a multivariate Gaussian distribution. Thus, $\max_{q,g,h}|S(q,g,h)|$ converges in distribution to the maximum of several multivariate Gaussian variables. We first estimate the covariance matrix of these statistics using their sample covariance. Then, we simulate random vectors from a Gaussian distribution with the same covariance matrix, take the absolute value of each component, and compute the maximum. By repeating this process, we obtain a series of bootstrap statistics. Finally, the threshold is set to the upper $\alpha$-quantile of these bootstrap statistics. If the test statistic exceeds this threshold, the null hypothesis is rejected. Theoretical properties of this test are detailed in Theorem 1 of Shi et al. (2020b) and Theorems 5 and 6 of Zhou et al. (2023b).